Race One Motors is an Indonesian car manufacturer. At its largest manufacturing facility, in Jakarta, the company produces subcomponents at a rate of 315 per day. and it uses these subcomponents at a -ate of 12, 330 per year (of 253 working days). Holding costs are $3 per item per year, and ordering costs are $31 per order.  


a) What is the economic production quantity?

b) How many production runs per year will be made? 
Answer #1

Demand = D = 12,300

Holding cost = Ch = $3

p = annual production capacity = per day production capacity*number of days = 315*250 = 78,750

Ordering cost = Co = $31

a. Economic production quantity = (2*D*Co/Ch)^1/2*(p/p-D)^1/2

= (2*12300*31/3)^1/2 * (78750/78750-12300)^1/2

= 504.18*1.09

= 548.87

b. Number of production runs = D/economic production quantity

= 12,300/548.87

= 22.41